Find the equation of a line that passes through the point (2,1) and has a gradient of

1

3

.

Leave your answer in the form

y

=

m

x

+

c


1
Expert's answer
2021-12-07T12:03:10-0500

for the line

"y=mx+c"

m is gradient

we have: "m=1\/3"

so, for point (2,1):

"1=2\/3+c"

"c=1-2\/3=1\/3"


equation of a line:

"y=x\/3+1\/3"